OpenStudy (anonymous):

Please...I stink at math What is the ratio 16 to 40 written as a fraction in lowest terms? A. 1/4 B. 8/20 C. 2/5 D. 3/8

geerky42 (geerky42):

What is 16 and 40 both divisible by?

geerky42 (geerky42):

Notice that both of them are even?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You can do this... 16/40=8/20=4/10=2/5

OpenStudy (anonymous):

I have 4 more...

OpenStudy (anonymous):

so c.

OpenStudy (anonymous):

Thank you. All of you.

OpenStudy (anonymous):

so c.